OpenStudy (anonymous):

If f(x)=x/3 and g(x) = 1/x find f(x)-g(x)

OpenStudy (agreene):

\[f(x)=\frac{x}3\] \[g(x)=\frac1{x}\] \[f(x)-g(x) = \frac{x}3-x^{-1}\] you can change the last bit if you want, depends on what youve been taught